# A simple fourth order propagator based on the Magnus expansion in the Liouville space: Application to a $\Lambda$-system and assessment of the rotating wave approximation

### v3

@@ I Introduction
Many quantum calculations involve time evolution of states or observables under time dependent Hamiltonian by nature or design. Well-known examples include nuclear magnetic resonance,[1, 2, 3, 4] quantum control (QC),[5, 6, 7, 8] the time-dependent Kohn-Sham equation,[9, 10] and more recently quantum computation[11, 12, 13, 14] and quantum sensing (QS).[15, 16, 17, 18] Recent advances in the formulation of driven open system quantum dynamics,[19, 20, 21, 22, 23] also brings the possibility that the time dependences are not only limited to Hamiltonians but can also be incorporated into the terms involving relaxation and decoherence. However, except for very few model systems and those with periodic time dependence, analytical solutions are not available when the system has general time dependence. For these cases, numerical solutions are inevitable.
@@
As the model system, we here consider a well-known $\Lambda$-system, which represents a simple three state system consisting of an excited state radiatively coupled to two low lying states that are not directly coupled with each other. The $\Lambda$-system has long served as a prototypical model for the description of laser driven quantum dynamical processes. Well known examples include QC,[5, 6, 43] coherent population trapping (CPT),[6, 44, 45] and electromagnetically induced transparency (EIT).[46, 47, 48, 49, 50, 51] More recently, $\Lambda$-system or extended versions have also been used as important model systems for QS,[15, 52, 53] for which more complicated pulses or time dependent fields are needed in general.
@@ II Numerical propagation based on Magnus expansion method
Considering the importance of open system dynamics under time-dependent Hamiltonian, either arising from an external driving or from noisy environments, extension of ME-based methods for open system dynamics have significant implications. To this end, we first review a simple fourth order method based on ME, which was derived recently,[29] and extend it to the application for an open system dynamics described by Lindblad equation.
@@ II.1 Unitary dynamics in the Hilbert space
In Eq. (2), the last term in the exponent involves a commutator, which introduces additional numerical errors. It is possible to remove this at the expense of increasing the number of exponentials, employing the Bakers-Campbell-Hausdorff formula.[41, 54] The resulting CFME version of Eq. (2) is
$\hat{U}^{(4)}_{{}_{\rm CF}}(t_{k+1},t_{k})=\exp\Bigl\{-\frac{i\delta t}{12\hbar}\left(3H_{k+1}+4H_{k+1/2}-H_{k}\right)\Bigl\}$
$\hskip 43.36243pt\times\exp\Bigl\{\frac{i\delta t}{12\hbar}\left(H_{k+1}-4H_{k+1/2}-3H_{k}\right)\Bigl\}.$ (3)
@@ II.2 Dynamics in the Liouville space
${\mathcal{D}}_{L}\hat{\rho}(t)=\sum_{j}c_{j}\left(\hat{L}_{j}\hat{\rho}(t)\hat{L}_{j}^{\dagger}-\frac{1}{2}\left\{\hat{L}_{j}\hat{L}_{j}^{\dagger},\hat{\rho}(t)\right\}_{+}\right),$ (5)
@@
For numerical propagation, it is convenient to propagate the density operator in the Liouville space, where it is viewed as a vector. The explicit form of ${\mathcal{L}}(t)$ in the Liouville space can be found from the corresponding definition in the Hilbert space by using the superket triple product identity,[56] as defined in Appendix A, and can be expressed as
@@
where the second equality defines the time-evolution superoperator $\mathcal{U}(t,0)$ in terms of $\mathcal{L}(\tau)$ given by Eq. (6). Appendix B also provides a proof that any finite truncation of ME for the Liouvillian in the Liouville space corresponds to a CPTP map.
For ${\mathcal{U}}(t,0)$, the extension of Eq. (2), namely our simplest 4th order propagator[29] based on ME is
@@
where ${\mathcal{L}}_{k}={\mathcal{L}}(t_{k})$. The above expression is equivalent to using Eq. (2) in the Hilbert space. The commutator free version of the above propagator has the same expression as Eq. (3) except that $H_{k}$ is replaced with ${\mathcal{L}}_{k}$.
@@
${\mathcal{U}}_{L}^{(4)}(t_{k+1},t_{k})=\exp\biggl\{-i\frac{\delta t}{6}\left(\mathcal{L}_{k}+4\mathcal{L}_{k+1/2}+\mathcal{L}_{k}\right)$
@@ III.2 Rotating wave approximation
Most of calculations and experimental interpretations involving the $\Lambda$-system were made under the rotating wave approximation (RWA), within which $\hat{H}_{mr}(t)$ is approximated as
@@
The RWA allows obtaining closed form expressions for eigenvalues of the Hamiltonian in the rotating wave frame (RWF), as detailed in Appendices C and D, and has been used widely. However, even for the simple $\Lambda$-system being considered here, to what extent the RWA holds and how it breaks down beyond weak or resonant fields are not clearly understood. It is widely accepted that the RWA is reasonable for weak and resonant condition. For the steady state case of $f_{p}(t)=f_{c}(t)=1$ considered here, this is easy to understand since the terms omitted in the above expression, namely, the counter-rotating terms, can be ignored as long as $\Omega_{\gamma}\ll\omega_{\gamma}$, where $\gamma=p$ or $c$, as follows:
@@
However, for intermediate and strong coupling strengths, the above condition breaks down and the RWA is expected to become inaccurate. For a quantitative assessment, we here provide detailed answers based on numerical calculations employing our ME-based propagators.
For the assessment of the RWA, we employ the following measure of error:
@@
where ${\hat{\rho}}^{\rm RWA}(t,0)$ is the system density operator evolving for the RWA Hamiltonian defined by Eq. (16) and the subscript $F$ denotes the Frobenius measure defined, for an operator $\hat{A}$, as follows: $|\hat{A}|_{F}=\sqrt{\sum_{i,j}|A_{ij}|^{2}}$, with $i$ and $j$ denoting the index of a matrix element of $\hat{A}$ in the basis of $|1\rangle$, $|2\rangle$, and $|3\rangle$.
@@ III.3 Numerical results for EIT and TPR regimes
We conducted numerical calculations for three cases of the parameters of the Hamiltonian. Case A corresponds to weak control and probe fields, whereas case B represents moderate control and probe fields. For case C, the control field is strong, breaking the condition for the RWA, whereas the probe field is moderate. For each case, we considered both closed unitary system dynamics without bath (I) and open system non-unitary dynamics (II) with couplings to photonic bath, for which the density operator evolves according to Eq. (4). The complete set of parameters are listed in Table LABEL:table-1. The units were chosen such that $\hbar=1$. Unless stated otherwise, we assume the two-photon resonance (TPR) condition with zero detuning[58] where $\omega_{p}=E_{2}-E_{1}$ and $\omega_{c}=E_{2}-E_{3}$. For these parameters, we chose $T=2\pi$. As the initial condition, we used $\hat{\rho}(0)=|1\rangle\langle 1|$.
@@
For the calculations in this subsection, we used our simplest 4th order propagators, Eqs. (2) and (11) in Sec. II, with a time step, $\delta t=T/(2^{10}-1)$. This choice was confirmed to give sufficiently accurate results when benchmarked against the results of our 6th order ME propagator[29] extended to the Liouville space. Thus, results presented in this subsection can be viewed virtually exact. Regarding our 6th order propagator, we would like to note that it is the original expression with commutators.[29] This is because a stable CFME for open system dynamics beyond 4th order was shown to be impossible.[59] Some of benchmark calculations are provided in the Supplementary Materials (SM). We also provide simplification of the algorithm that allows efficient long time dynamics for periodic Hamiltonian in App. E, which can be used for arbitrarily long time.
@@
Figure 2 shows errors of RWA calculated according to Eq. (18) for all six cases during the initial dynamics. As expected, the errors for cases A-I and A-II, which correspond to weak control and probe fields, are smallest of the three cases and are negligible. The errors for cases B-I and B-II are about an order of magnitude larger and quantitatively significant although small. It is interesting to note that errors for the open system dynamics (B-II) are about half of those for closed system dynamics (B-I). On the other hand, the errors for the cases C-I and C-II grow rapidly during the initial stage and remain substantial. The presence of the bath for this case (C-II) is also shown to reduce the overall error but appears to be less effective than the case B-II. These are expected results since the strong control field renders the RWA to be inaccurate even though the probe field is moderate. Nonetheless, quantitative details provided here can serve as a meaningful reference for assessing the applicability of RWA. The SM shows errors in the long time limits for the three open system quantum dynamics, which confirm these assessments in the presence of bath.
@@
Figure 3: Elements of the time dependent system density operators for the case B-I, closed system unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the 4th order ME-propagator with commutator, Eq. (8).
@@
Figure 4: Elements of the time dependent system density operator for the case B-II, the open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA (dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the order ME-propagator with commutator, Eq. (11).
@@
As a representative example, for the cases B-I and B-II, we here provide detailed comparisons of time evolving matrix elements of the density operators between the full and RWA Hamiltonians. The results for A-I and A-II, for which the RWA is expected to work well, and those for cases C-I and C-II, for which the RWA is expected to clearly break down, are provided in Figs. S1-S4 in the SM.
@@
Figure 3 shows elements of time evolving density operator for the case B-I in Table LABEL:table-1, which corresponds to a representative example of EIT with perfect resonance but without the effects of bath. Results for both the full Hamiltonian and the RWA are shown. We use $\tilde{\rho}_{ij}$ to represent the off-diagonal elements in the RWF (see Appendix C). As expected, the two results are in reasonable agreement. In particular, all the diagonal elements of the density operator based on the RWA reproduce those of exact dynamics, except for small wiggles that originate from non-resonant terms of the Hamiltonian. The errors in off-diagonal elements of the density operator are reasonably small but are slightly worse than diagonal elements. In particular, the real parts of the off-diagonal elements of the density operator are shown to exhibit larger deviation.
@@
Figure 4 shows results for the case B-II in Table LABEL:table-1, for which EIT is observed with full resonance condition in the presence of photonic bath. The density operator in this case is governed by Eq. (4). Overall, the presence of bath significantly dampens coherent behavior and reduces the discrepancies between the results for the full and RWA Hamiltonians. It is interesting to note that both real and imaginary parts of $\tilde{\rho}_{23}$ for this open system dynamics are smaller than those for the closed system dynamics by an order of magnitude, whereas the errors of the RWA from the exact dynamics for these off-diagonal terms remain of similar order. As a result, the relative error of RWA for these off-diagonal elements of the density operator turn out to be larger than other elements.
@@
In order to understand implications of the errors due to the RWA for the EIT behavior, we conducted calculations for the long time limit for values of $\omega_{p}$ from 2 to 10 in increments of 0.25. All of these cases correspond to $T=8\pi$. Once we have conducted the long time dynamics, we then calculated the average steady-state values of the system density operator matrix elements by integrating their time dependent values over an interval of $8\pi$ with the trapezoid rule. For the case A-II, Fig. 5 shows that most results for RWA agree very well with those for the full Hamiltonian. The results for the case B-II are shown in Fig. 6. All the diagonal and off-diagonal elements based on the RWA are in reasonable agreement with exact values over the range of $\delta\omega_{p}$ tested. On the other hand, Fig. 7 shows significant qualitative differences between the two density operators for the case C-II, which confirms the unreliability of the RWA for the case of strong control fields.
@@
Figure 9: Steady state limits of populations at state 2 ($\rho_{22}$) for the open system of case C-II with full Hamiltonian (a) and RWA (b) the RWA, for which times required for convergence are respectively plotted in (c) and (d).
Note that results presented so far have not covered some interesting parameter regimes, such as the TPR with a non-zero detuning ($\delta\omega_{p}=\delta\omega_{c}\neq 0$), for which the phenomenon of CPT can still be observed. In recent works,[60, 61] these modified detunings were used in order to avoid transitions to an unwanted level, [60] and to swap populations of the two lower-lying states.[61] Figures 8 and 9 provide results for Cases B-II and C-II, respectively, in this regime and other parameter regimes by using $\omega_{c}\ \in[1,1.25,...,6.75,7]$ and $\omega_{p}\ \in[2,2.25...,9.75,10]$. A similar figure for Case A-II is provided in the SI.
For weak and moderate field strengths, it is seen that there is no population in the excited state in the steady state limit under TPR conditions (the main diagonal in panels (a) and (b)). This explains why TPR is a requirement for stimulated Raman adiabatic passage (STIRAP),[62] which aims to transfer population from state $|1\rangle$ to state $|3\rangle$ without significantly populating the excited state. However, for the strong control field case of C-II in Fig. 9, we see that the RWA population in panel (b) is a poor predictor of the true population in panel (a). In particular, the population is a much more complicated function of detunings, and the intermediate state 2 has non-negligible population even when the TPR condition under the assumption of the RWA is satisfied. This suggests principles of STIRAP need to be examined more carefully beyond the RWA.

### v4

@@ I Introduction
Many quantum calculations involve time evolution of states or observables under time dependent Hamiltonian by nature or design. Well-known examples include nuclear magnetic resonance,[1, 2, 3, 4] quantum control (QC),[5, 6, 7, 8, 9, 10] quantum thermodynamics,[11, 12, 13, 14] the time-dependent Kohn-Sham equation,[15, 16] and more recently quantum computation[17, 18, 19, 20] and quantum sensing (QS).[21, 22, 23, 24] Recent advances in the formulation of driven open system quantum dynamics,[25, 14, 26, 27, 28, 29, 30] also brings the possibility that the time dependences are not only limited to Hamiltonians but can also be incorporated into the terms involving relaxation and decoherence. However, except for very few model systems and those with periodic time dependence, analytical solutions are not available when the system has general time dependence. For these cases, numerical solutions are inevitable.
@@
As the model system, we here consider a well-known $\Lambda$-system, which represents a simple three state system consisting of an excited state radiatively coupled to two low lying states that are not directly coupled with each other. The $\Lambda$-system has long served as a prototypical model for the description of laser driven quantum dynamical processes. Well known examples include QC,[5, 6, 50] coherent population trapping (CPT),[6, 51, 52] and electromagnetically induced transparency (EIT).[53, 54, 55, 56, 57, 58] More recently, $\Lambda$-system or extended versions have also been used as important model systems for QS,[21, 59, 60] for which more complicated pulses or time dependent fields are needed in general. Many of these studies have relied on the rotating wave approximation (RWA), which has been well tested for weak matter-radiation interaction and resonant conditions. However, beyond these well-established regimes, the errors of RWA are poorly understood even for the $\Lambda$-system. Thus, we provide an assessment of RWA in this work.
@@ II Numerical propagation based on Magnus expansion method
Considering the importance of open system dynamics under time-dependent Hamiltonian, either arising from an external driving or from noisy environments, extension of ME-based methods for open system dynamics have significant implications. To this end, we first review a simple fourth order method based on ME, which was derived recently,[36] and extend it to the application for an open system dynamics described by the Lindblad equation.[41]
@@ II.1 Unitary dynamics in the Hilbert space
In Eq. (2), the last term in the exponent involves a commutator, which introduces additional numerical errors. It is possible to remove this at the expense of increasing the number of exponentials, employing the Baker-Campbell-Hausdorff formula,[48, 49] as described in App. A. The resulting CFME version of Eq. (2) is
$\hat{U}^{(4)}_{{}_{\rm CF}}(t_{k+1},t_{k})=\exp\Bigl\{-\frac{i\delta t}{12\hbar}\left(3\hat{H}_{k+1}+4\hat{H}_{k+1/2}-\hat{H}_{k}\right)\Bigl\}$
$\hskip 43.36243pt\times\exp\Bigl\{\frac{i\delta t}{12\hbar}\left(\hat{H}_{k+1}-4\hat{H}_{k+1/2}-3\hat{H}_{k}\right)\Bigl\}.$ (3)
@@ II.2 Dynamics in the Liouville space
${\mathcal{D}}_{L}\hat{\rho}(t)=\sum_{j}c_{j}\left(\hat{L}_{j}\hat{\rho}(t)\hat{L}_{j}^{\dagger}-\frac{1}{2}\left\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\hat{\rho}(t)\right\}_{+}\right),$ (5)
@@
For numerical propagation, it is convenient to propagate the density operator in the Liouville space, where it is viewed as a vector. The explicit form of ${\mathcal{L}}(t)$ in the Liouville space can be found from the corresponding definition in the Hilbert space by using the superket triple product identity,[62] as defined in App. B, and can be expressed as
@@
where the second equality defines the time-evolution superoperator $\mathcal{U}(t,0)$ in terms of $\mathcal{L}(\tau)$ given by Eq. (6). Appendix C also provides a proof that any finite truncation of ME for the Liouvillian in the Liouville space corresponds to a CPTP map.
For ${\mathcal{U}}(t,0)$, the extension of Eq. (2), namely our simplest 4th order propagator[36] based on ME, is
@@
where ${\mathcal{L}}_{k}={\mathcal{L}}(t_{k})$. The above expression is equivalent to using Eq. (2) in the Hilbert space. The commutator free version of the above propagator has the same expression as Eq. (3) except that $\hat{H}_{k}$ is replaced with ${\mathcal{L}}_{k}$.
@@
${\mathcal{U}}_{L}^{(4)}(t_{k+1},t_{k})=\exp\biggl\{-i\frac{\delta t}{6}\left(\mathcal{L}_{k}+4\mathcal{L}_{k+1/2}+\mathcal{L}_{k+1}\right)$
@@ III.2 Rotating wave approximation
Most of calculations and experimental interpretations involving the $\Lambda$-system were made under the RWA, within which $\hat{H}_{mr}(t)$ is approximated as
@@
The RWA allows obtaining closed form expressions for eigenvalues of the Hamiltonian in the rotating wave frame (RWF), as detailed in Appendixes D and E, and has been used widely. It is widely accepted that the RWA is reasonable for weak and resonant condition. For the steady state case of $f_{p}(t)=f_{c}(t)=1$ considered here, this is easy to understand since the terms omitted in the above expression, namely, the counter-rotating terms, can be ignored as long as $\Omega_{\gamma}\ll\omega_{\gamma}$, where $\gamma=p$ or $c$, as follows:
@@
However, for intermediate and strong coupling strengths, the above condition breaks down and the RWA is expected to become inaccurate. While this is expected, quantitative details of how such breakdown occurs are not clearly known even for the simple $\Lambda$-system considered here. Thus, we here provide such an assessment employing our ME-based propagators.
For a quantitative assessment of the accuracy of the RWA, we employ the following measure of error:
@@
where ${\hat{\rho}}^{\rm RWA}(t)$ is the system density operator that evolves according to the RWA Hamiltonian, Eq. (16). The subscript $F$ in Eq. (18) denotes the Frobenius measure , defined as $|\hat{A}|_{F}=\sqrt{\sum_{i,j}|A_{ij}|^{2}}$ for an operator $\hat{A}$, with $i$ and $j$ indicating the matrix element in the basis of $|1\rangle$, $|2\rangle$, and $|3\rangle$.
@@ III.3 Numerical results for EIT and TPR regimes
We conducted numerical calculations for three cases of the parameters of the Hamiltonian. Case A corresponds to weak control and probe fields, whereas case B represents moderate control and probe fields. For case C, the control field is strong, breaking the condition for the RWA, whereas the probe field is moderate. For each case, we considered both closed unitary system dynamics without bath (I) and open system non-unitary dynamics (II) with couplings to photonic bath, for which the density operator evolves according to Eq. (4). The complete set of parameters are listed in Table LABEL:table-1. The units were chosen such that $\hbar=1$. Unless stated otherwise, we assume the two-photon resonance (TPR) condition with zero detuning[64] where $\omega_{p}=E_{2}-E_{1}$ and $\omega_{c}=E_{2}-E_{3}$. For the initial condition, we chose $\hat{\rho}(0)=|1\rangle\langle 1|$.
If the Hamiltonian is periodic in time with period $T$, explicit calculation of the propagator up to $t=T$ is sufficient to generate the dynamics for arbitrarily long time, as described in App. F. Indeed, for all the parameters of Table LABEL:table-1 and under the zero-detuning TPR condition, the Hamiltonian is periodic in time, with $T=2\pi$. Thus, we use the approach in App. F to generate the dynamics for $t>T$ including the steady state limits. This way, we can calculate accurate enough steady state limit results without resorting to Floquet approaches.[65, 25, 66, 67, 68, 69, 70]
@@
For all the calculations in this subsection, we used our simplest 4th order propagators, Eqs. (8) and (11) in Sec. II, with a time step, $\delta t=T/(2^{10}-1)$. This choice was confirmed to give sufficiently accurate results when benchmarked against the results of our 6th order ME propagator[36] extended to the Liouville space. Thus, results presented in this subsection can be viewed virtually exact. Regarding our 6th order propagator, we would like to note that it is the original expression with commutators.[36] This is because a stable CFME for open system dynamics beyond 4th order with real coefficients was shown to be impossible.[71] Some of benchmark calculations are provided in the Supplementary Materials (SM).
@@
Figure 2 shows errors of the RWA calculated according to Eq. (18) for all six cases during the initial dynamics. As expected, the errors for cases A-I and A-II, which correspond to weak control and probe fields, are smallest of the three cases and are negligible. Cases B-I and B-II exhibit about an order of magnitude larger errors although small. It is interesting to note that errors for the open system dynamics (B-II) are about half of those for closed system dynamics (B-I). On the other hand, the errors for the cases C-I and C-II grow rapidly during the initial stage and remain substantial. The presence of the bath for this case (C-II) is also shown to reduce the overall error, although less effective than the case B-II. These are expected results since the strong control field renders the RWA to be inaccurate even though the probe field is moderate. Nonetheless, quantitative details provided here offer a meaningful reference for assessing the applicability of RWA. The supplementary material shows errors in the long time limits for the three open system quantum dynamics, which confirm these assessments in the presence of bath.
@@
Figure 3: Elements of the time dependent system density operators for the case B-I, closed system unitary dynamics, with full Hamiltonian (solid) and the RWA(dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the fourth-order ME-propagator with commutator, Eq. (8).
@@
Figure 4: Elements of the time dependent system density operator for the case B-II, the open system non-unitary dynamics, with full Hamiltonian (solid) and the RWA (dashed), for which the Hamiltonians are respectively $\hat{H}(t)$ and $\hat{H}^{\rm\small RWA}(t)$. Both were calculated using the fourth-order ME-propagator with commutator, Eq. (11).
@@
As a representative example, we next provide further details of the difference in density operators for the full and RWA Hamiltonians for cases B-I and B-II. The results for cases A-I and A-II, where the RWA works well, and cases C-I and C-II, where the RWA clearly breaks down, are provided in Figs. S1-S4 in the SM.
@@
Figure 3 shows elements of time evolving density operator for the case B-I in Table LABEL:table-1, which corresponds to a representative example of EIT with perfect resonance but without the effects of bath. We use $\tilde{\rho}_{ij}$ to represent the off-diagonal elements in the RWF (see App. D). Overall, the results for the RWA Hamiltonian are in reasonable agreement with those for the full Hamiltonian. In particular, all the diagonal elements of the density operator based on the RWA reproduce those of of the full Hamiltonian, except for small wiggles that originate from non-resonant terms of the Hamiltonian. The errors in off-diagonal elements of the density operator are small but slightly worse than those for diagonal elements. In particular, the real parts of the off-diagonal elements of the density operator are shown to exhibit larger deviation.
@@
Figure 4 shows results for the case B-II in Table LABEL:table-1, for which EIT is observed with full resonance condition in the presence of photonic bath. The time evolution of the density operator in this case is governed by Eq. (4). Overall, the presence of bath significantly dampens coherent behavior and reduces the discrepancies between the results for the full and RWA Hamiltonians. It is interesting to note that both real and imaginary parts of $\tilde{\rho}_{23}$ for this open system dynamics are smaller than those for the closed system dynamics by an order of magnitude, whereas the errors of the RWA from the exact dynamics for these off-diagonal terms remain of similar order. As a result, the relative error of RWA for these off-diagonal elements of the density operator turn out to be larger than other elements.
@@
In order to understand implications of the errors due to the RWA for the EIT behavior, we conducted calculations for the long time limit for values of $\omega_{p}$ from 2 to 10 in increments of 0.25. For all of these cases, the Hamiltonian remain periodic in time with a common period, $T=8\pi$. Thus, the approach of App. F can still be used to effectively calculate the long time propagation. Once we have conducted the long time dynamics, we then calculated the average steady-state values of the system density operator matrix elements by integrating their time dependent values over an interval of $8\pi$ with the trapezoid rule. For the case A-II, Fig. 5 shows that most results for RWA agree very well with those for the full Hamiltonian. The results for the case B-II are shown in Fig. 6. All the diagonal and off-diagonal elements based on the RWA are in reasonable agreement with exact values over the range of $\delta\omega_{p}$ tested. On the other hand, Fig. 7 shows that there is significant qualitative difference between the two density operators for the case C-II, which confirms the unreliability of the RWA for the case of strong control fields.
@@
Figure 9: Steady state limits of populations at state 2 ($\rho_{22}$) for the open system of case C-II with full Hamiltonian (a) and RWA (b), for which times required for convergence are respectively plotted in (c) and (d).
Note that results presented so far do not cover the parameter regime where the TPR with a non-zero detuning ($\delta\omega_{p}=\delta\omega_{c}\neq 0$) occurs, for which the phenomenon of CPT can still be observed. In recent works,[72, 73] these modified detunings were used in order to avoid transitions to an unwanted level, [72] and to swap populations of the two lower-lying states.[73] Figures 8 and 9 provide such results for cases B-II and C-II, respectively, for $\omega_{c}\ \in[1,1.25,...,6.75,7]$ and $\omega_{p}\ \in[2,2.25...,9.75,10]$. A similar figure for case A-II is provided in the SM.
For weak and moderate field strengths, it is seen that there is no population in the excited state in the steady state limit under TPR conditions (the main diagonal in panels (a) and (b)). This explains why TPR is a requirement for stimulated Raman adiabatic passage (STIRAP),[74] which aims to transfer population from state $|1\rangle$ to state $|3\rangle$ without significantly populating the excited state. However, for the strong control field case of C-II in Fig. 9, we see that the RWA population in panel (b) is a poor predictor of the true population in panel (a). In particular, the population is a much more complicated function of detunings, and the intermediate state 2 has non-negligible population even when the TPR condition under the assumption of the RWA is satisfied. This suggests that principles of STIRAP need to be examined more carefully beyond the RWA.